# Partially Retargeted Balancing Weights for Causal Effect Estimation Under Positivity Violations

### v1

@@ Abstract
Positivity violations pose significant challenges for causal effect estimation with observational data. Under positivity violations, available methods result in either treatment effect estimators with substantial statistical bias and variance or estimators corresponding to a modified estimand and target population that is misaligned with the original research question. To address these challenges, we propose partially retargeted balancing weights, which yield reduced estimator variance under positivity violations by modifying the target population for only a subset of covariates. Our weights can be derived under a novel relaxed positivity assumption allowing the calculation of valid balancing weights even when positivity does not hold. Our proposed weighted estimator is consistent for the original target estimand when either 1) the implied propensity score model is correct; or 2) the subset of covariates whose population is not modified contains all treatment effect modifiers. When these conditions do not hold, our estimator is consistent for a slightly modified treatment effect estimand. Furthermore, our proposed weighted estimator has reduced asymptotic variance when positivity does not hold. We evaluate our weights and corresponding estimator through applications to synthetic data, an EHR study, and when transporting an RCT treatment effect to a Midwestern population.
@@ Introduction
Confounding is a key challenge when estimating causal effects of binary treatments with observational data. While many methods exist for controlling confounding by balancing covariates between treatment groups, these methods do not fully control confounding when there are positivity violations. Positivity violations occur when there exists at least one combination of covariate values which perfectly predicts treatment assignment, a phenomenon which is especially likely to occur with high-dimensional covariates (D’Amour and Franks,, 2021). When there are positivity violations, inverse probability weighting (IPW) estimators (Rosenbaum and Rubin,, 1983; Hahn,, 1998; Robins et al.,, 2000; Hirano and Imbens,, 2001; Hirano et al.,, 2003; Imbens,, 2004) have substantial statistical bias and inflated variance due to extreme weights. A proposed alternative to IPW, direct balancing weights, (Hainmueller,, 2012; Zubizarreta,, 2015; Wang and Zubizarreta,, 2020) tend to yield reduced error and variance compared to IPW methods through achieving exact covariate balance for a prespecified set of covariates. However, when there are positivity violations there is often no solution to the direct balancing optimization problem prohibiting their use altogether.
@@
To mitigate the impacts of positivity violations on causal effect estimation, a variety of methods, such as propensity score trimming, matching weights, and overlap weights, have been developed that modify the estimand to one whose target population has increased overlap to achieve estimators with reduced variance (Crump et al.,, 2009; Li and Greene,, 2013; Yang and Ding,, 2018; Li et al.,, 2018, 2019). In fact, overlap weights (Li et al.,, 2018) do not require positivity at all by targeting the so-called ‘overlap population’; this result in an estimator with minimum asymptotic variance among weighted treatment effects under homoscedasticity. However, these methods target a population that is often substantially different than the original population of interest such that the corresponding treatment effect estimate may be misaligned with the original research question. Yet, a key benefit of causal effect estimation with observational data is the ability to evaluate causal effects in a meaningful population of interest (e.g., transporting effects) such that modifying the estimand is either undesirable or challenging to interpret. Furthermore, the corresponding estimators are biased for the original estimand when there is treatment effect heterogeneity. Modified treatment policy estimands result in improved estimator performance by defining an entirely different estimand arrived at by imagining modifications to treatment (Kennedy,, 2019). While these estimators tend to perform well, these modified treatments tend to be more challenging to conceptualize than binary treatments.
@@
Thus, under positivity violations, methods either 1) target an estimand of interest but result in poor estimator performance or technical issues (e.g., balancing weights); or 2) alleviate these technical issues by targeting a modified estimand that may be misaligned with the original estimand of interest (e.g., overlap weights). To address these challenges of estimator variance, bias, and interpretation, we propose a novel direct balancing approach whose optimization problem yields a solution under positivity violations by modifying the target population for only a subset of the covariates. In doing so, our proposed weights achieve exact covariate balance between treated and control groups for all covariates; in contrast, minimal weights (Wang and Zubizarreta,, 2020) achieve a solution through by focusing on approximate, rather than exact, covariate balance. The target population under our approach can be modified as minimally as possible to achieve a solution and the corresponding estimator tends to minimize bias and variance with respect to the modified estimand. When the target population is instead modified for all covariates, overlap weights are a special case of our proposed weights. Our proposed estimator is consistent for the original estimand of interest when the either the implied propensity score model is correct or the subset of covariates whose population is modified are not treatment effect modifiers. We derive asymptotic results which show that 1) our proposed approach explicitly relaxes the positivity assumption; and 2) our proposed estimator has reduced asymptotic variance when there are positivity violations. We propose a design- and model-based procedure for selecting the subset of the covariates not balanced to the target population and we evaluate the corresponding estimators through applications to synthetic data, the MIMIC-III study, and when transporting the effect of a healthcare hot spotting study.
@@
The remainder of the paper is organized as follows. In Section 2, we introduce notation and methodological background on direct balancing and minimal weights. We introduce our proposed partially retargeted direct balancing approach and explore the corresponding relaxed positivity assumption, interpretation of the implied propensity score, and outcome models in Section 3. In Section 4, we discuss the asymptotic properties of our proposed estimator and in Section 5 we discuss adaptations of proposed method to other estimation problems. We discuss implementation details and propose design- and model-based estimators for our method in Section 6. In Sections 7, 8, and 9, we apply our methods to synthetic data, an EHR study, and transporting a health care hotspotting RCT effect.
@@
Consider an independent sample $\{(Z_{i},Y_{i},\bm{X}_{i})\}_{i=1}^{n}$ of size $n$ from a population where $Z_{i}=z$, indicates belonging to the treatment ($z=1$) or control group ($z=0$), $Y_{i}$ is the outcome of interest, and $\bm{X}_{i}=(X_{i1},\ldots X_{ip})$ is a $p$ length vector of pre-treatment covariates. Let $Y(z)$ be the potential outcome that would be observed if assigned to treatment group $z$ (Neyman, (1990); Rubin, (1974, 1978); Hernan and Robins, (2024)). We assume the standard stable unit treatment value assumption (SUTVA) which implies that $Y_{i}=Y_{i}(Z_{i})$ such that we only observe one potential outcome for each individual. We focus on the average treatment effect (ATE), $\tau=E[Y(1)-Y(0)]$ as the estimand of interest, although this work can be readily extended to other estimands of interest which we discuss in Section 5. Let $e(\bm{x})=\Pr(Z_{i}=1|\bm{X}_{i}=\bm{x})$ be the propensity score. For the identification of $\tau$ the following assumptions are typically made:
@@
The focus of our work surrounds the positivity assumption, specifically, scenarios where this assumption does not hold. Let $\mu_{z}(\bm{x})=E[Y(z)|\bm{X}=\bm{x}]$ and $\tau(\bm{x})=\mu_{1}(\bm{x})-\mu_{0}(\bm{x})$. Then the following identification holds with the above assumptions:
$\tau$ $=E\left\{\frac{\mu_{1}(\bm{X})Z}{e(\bm{X})}-\frac{\mu_{0}(\bm{X})(1-Z)}{1-e(\bm{X})}\right\}=E\left\{\frac{Y(1)Z}{e(\bm{X})}-\frac{Y(0)(1-Z)}{1-e(\bm{X})}\right\}=E\left\{\frac{YZ}{e(\bm{X})}-\frac{Y(1-Z)}{1-e(\bm{X})}\right\},$
@@
where the sample version of the last equation corresponds to the standard inverse probability weighting (IPW) estimator of the ATE. For a treatment effect estimator be unbiased, the expectations $\mu_{0}(\bm{x})$ and $\mu_{1}(\bm{x})$ need to be balanced between the treated and control populations and to the target population. IPW balance all functions of $\bm{x}$ in this way such that the following equations hold,
$E\left\{\frac{Zb(\bm{X})}{e(\bm{X})}\right\}=E\{b(\bm{X})\}\>\>$ $\text{and}\>\>E\left\{\frac{(1-Z)b(\bm{X})}{1-e(\bm{X})}\right\}=E\{b(\bm{X})\}$ (1)
$E\left\{\frac{Zb(\bm{X})}{e(\bm{X})}\right\}$ $=E\left\{\frac{(1-Z)b(\bm{X})}{1-e(\bm{X})}\right\}$ (2)
@@
for any function $b(\bm{x})$, which we reference as three-way balance (Chan et al.,, 2016). Thus, IPW satisfy three-way balance of both $\mu_{0}(\bm{x})$ and $\mu_{1}(\bm{x})$, regardless of their true forms. Specifically, Equation (1) balances the weighted treated and control group expectations of $b(\bm{x})$ to the population expectation of $b(\bm{x})$ and Equation (2) balances the weighted treated and control group expectations of $b(\bm{x})$ to each other. Thus, Equation (1) and (2) serve different purposes; while Equation (1) implies Equation (2), Equation (1) primarily serves to balance $b(\bm{x})$ in the treated and control groups to $b(\bm{x})$ to the estimand target population. In contrast, Equation (2) only ensures that $b(\bm{x})$ is balanced between treated and control groups. When only Equation (2) holds, the corresponding treatment effect estimator will be consistent for a modified target population and causal estimand but could not be consistent for the ATE.
@@
However, when there is a lack of overlap, even approximate balance of the empirical counterparts of Equations (1) and (2) may not be achieved after IPW, resulting in estimators with substantial bias and variance. To achieve balance when there is a lack overlap, some methods target a different estimand, $E[h(\bm{X})\tau(\bm{X})]$, which corresponds to a modified target population with improved overlap. For this modified estimand, Equations (1) and (2) must hold for $h(\bm{x})b(\bm{x})$. Li et al., (2018) propose one such method, overlap weights ($h(\bm{x})=e(\bm{x})(1-e(\bm{x})$), which are specifically designed to ensure that the weights exist even when positivity does not hold. By modifying the target population, overlap weights can more easily satisfy Equation (1) yielding an estimator with reduced bias and variance with respect to the modified estimand, the ATO. However, the overlap weights estimator tends to be biased for the ATE if there is treatment effect heterogeneity and the modified population is often challenging to characterize with respect to the original research question. Propensity score trimming (Crump et al.,, 2009) also modifies the target population to achieve improved overlap by discarding observations.
While the weights corresponding to these modified estimands tend to achieve improved finite-sample balance in scenarios with a lack of overlap, they rarely achieve exact finite-sample balance of empirical versions of Equations (1) and (2), which results in error in the corresponding treatment effect estimator. In fact, even when positivity does hold, exact finite-sample balance is rarely achieved by IPW-style estimators. While exact balance of covariate sample means is achieved for the ATO when the propensity score is estimated with a logistic regression (Li et al.,, 2018), for a majority of estimands, exact balance of the sample means will not be achieved, especially when there are positivity violations. This has motivated the development of sample weights that explicitly achieve three-way finite sample balance for a set of covariate functions.
@@
$\frac{1}{n}\sum_{i=1}^{n}w_{i}I[Z_{i}=z]b_{j}(\bm{X}_{i})-\frac{1}{n}\sum_{i=1}^{n}b_{j}(\bm{X}_{i})$ $=0\>\>\>j=1,\ldots,J,\>\>\>z=0,1$ (3b)
where $D(\cdot)$ is a convex measure of dispersion. The constraints $w_{i}\geq 0$ and $\sum_{i=1}^{n}w_{i}Z_{i}=\sum_{i=1}^{n}w_{i}(1-Z_{i})$ are often added to ensure that the weights do not extrapolate the observed outcomes. The resulting direct balancing weights estimator is
@@
In this constrained optimization problem, constraint (3b) ensures that the treated and control covariate functions are balanced to the sample population. These constrains also imply that the treated and control covariate functions are balanced to each other (i.e., $\frac{1}{n}\sum_{i=1}^{n}w_{i}Z_{i}b_{j}(\bm{X}_{i})=\frac{1}{n}\sum_{i=1}^{n}w_{i}(1-Z_{i})b_{j}(\bm{X}_{i})$) such that three-way balance is satisfied by $\{w^{bw}_{i}\}_{i=1}^{n}$ for $b(\bm{x})$. Since the treated and control balance is implicitly satisfied, Problem (3) is separable by treatment group such that two separate constrained optimization problems can be solved to derive $\{w^{bw}_{i}\}_{i=1}^{n}$ (Chan et al.,, 2016); we discuss implications of this separation in Section 3.1.
When there is a lack of overlap between $\{b(\bm{x})\}_{z=1}$ and $\{b(\bm{x})\}_{z=0}$, often there is no solution to Problem (3) (i.e., Equations (3a)-(3b)) such that weights which achieve exact three-way balance cannot be derived. In these scenarios, the exact balance constraint is relaxed and approximate balance is achieved by constraining the absolute difference between sample means to be less than or equal to some parameter $\delta$; weights derived through this process are called minimal weights Wang and Zubizarreta, (2020). While the minimal weights estimator does have reduced variance compared to IPW and direct balancing estimators, minimal weights achieve none of the three balancing equations. Therefore, the minimal weights estimator may have bias due to both deviation from the target population and covariate imbalance. Thus, there is a need for methods that 1) achieve exact finite-sample covariate balance between treated and control distributions to reduce estimator error; and 2) modify the target population minimally and/or in an interpretable manner to yield an estimator with reduced variance when there is a lack of overlap.
A new partially retargeted balancing approach under positivity violations
To address these challenges, we propose a novel constrained optimization problem for deriving weights, primarily for scenarios where there is no solution to direct balancing weights due to positivity violations. Consider the case where Assumption Assumption 2: Positivity does not hold, but $0<\text{Pr}\{Z_{i}=1|c(\bm{X})=c(\bm{x})\}$ for some $c(\bm{x})\subset b(\bm{x})$ such that positivity holds when conditioning on this restricted subset. Let $g(\bm{x})$ be the remaining covariate functions in $b(\bm{x})$ such that $b(\bm{x})=[c(\bm{x})=\{c_{k}(\bm{x})\}_{k=1}^{K},g(\bm{x})=\{g_{l}(\bm{x})\}_{l=1}^{L}]$. The key idea of our method is to relax the constraints by balancing all covariate functions, $b(\bm{x})=\{c(\bm{x}),g(\bm{x})\}$, between treated and control groups, but only $c(\bm{x})$ to the sample population, yielding the following set of balancing equations:
@@
The primary difference between this constrained optimization problem and the direct balancing weights constrained optimization problem is Equation (5c); in this equation, the functions, $g(\bm{x})=\{g_{l}(\bm{x})\}_{l=1}^{L}$ are only balanced between treated and control groups and not to the sample population. Thus, the derived weights satisfy exact three-way balance for functions $c(\bm{x})=\{c_{k}(\bm{x})\}_{k=1}^{K}$ but only balance between treated and control distributions for $g(\bm{x})$. In doing so, the derived weights do not balance $g(\bm{x})$ to a specific modified population (such as with overlap weights), however, the weighted treated group sample means of $g(\bm{x})$ are implicitly modified from the sample population mean when minimizing weight dispersion subject to only the treated and control balance constraint. We call these weights, $\{w_{i}^{*}\}_{i=1}^{n}$, partially retargeted balancing weights (PRTBW). In Section 3.1, we formally show that this approach is justified under a significantly relaxed positivity assumption. In addition, the corresponding estimator to Problem (5) (i.e., Equations (5a) - (5c)) is consistent for the ATE when $c(\bm{x})$ contains all treatment effect modifiers, which we discuss further in Section 3.2.
@@
Figure 1: Simulated data exhibiting the different between direct balancing weights, minimal weights, and the proposed partially retargeted balancing weights. Squares (purple) and triangles (orange) indicate the unweighted control and treated sample means, respectively. The filled diamond indicates the sample mean while the outlined diamond indicates the weighted sample mean given $\{w_{i}^{*}\}$ derived through Problem (5). In the right-hand figures, the horizontal line indicates the sample mean while the vertical lines end at the weighted treated and control means for balancing weights (BW), minimal weights (MW) and our proposed weights (PRTBW). For this simulated data, there is no solution to Problem (3), exhibited by the dashed vertical line.
@@
Figure 1 presents a toy example of the differences in covariate balance and target population between balancing weights (Problem (3)), minimal weights, and our proposed partially retargeted balancing weights (Problem (5)). For this simulated data, there is no solution to Problem (3), such that weights that exactly balance the treated and control sample means of $b(\bm{x})=\{c(\bm{x}),g(\bm{x})\}$ cannot be derived. Our proposed weights balance the treated and control sample means of $c(\bm{x})$ to the sample mean, but the treated and control sample means of $g(\bm{x})$ are balanced to a value that is substantially above the sample mean, providing a viable solution to Problem (5). In contrast, minimal weights do not achieve covariate balance for either $c(\bm{x})$ or $g(\bm{x})$. In fact, the weighted treated sample means are unchanged and for both $c(\bm{x})$ and $g(\bm{x})$ and the weighted control sample mean of $c(\bm{x})$ is further away from the sample mean of $c(\bm{x})$ than the unweighted control sample mean. Thus, while both the minimal weights optimization problem and our proposed Problem (5) yield solutions for this data, they do this through different mechanisms and each yields a solution with a different interpretation.

### v2

@@ Abstract
Positivity violations, which occur when some subgroups either always or never receive a treatment of interest, pose significant challenges for causal effect estimation with observational data. Recent balancing weight methods have proved to be highly effective in confounding control, however their utility is diminished in the presence of positivity violations, resulting in bias and excess variance. Approaches that deal with positivity violations, on the other hand, work by targeting a modified estimand that may be misaligned with the original research question. To address these challenges, we propose a novel balancing weights approach, which mitigates positivity violations while attempting to retain the original estimand by a targeted relaxation of the balancing constraints. Our proposed weighted estimator is consistent for the original estimand when either 1) the implied propensity score model is correct; or 2) all treatment effect modifiers are balanced to the target population. When these conditions do not hold, our estimator is consistent for a slightly modified treatment effect estimand. Furthermore, our proposed weighted estimator has reduced asymptotic variance when positivity does not hold. We evaluate our approach through applications to synthetic data, an observational study, and when transporting a treatment effect from a randomized trial.
@@ Introduction
Confounding is a key challenge when estimating the causal effect of a binary treatment with observational data. While many methods exist for controlling confounding by balancing covariates between treatment groups, these methods do not fully control confounding when there are positivity violations. Positivity violations occur when there exists at least one combination of covariate values which perfectly predicts treatment assignment, a phenomenon which is especially likely to occur with high-dimensional covariates (D’Amour and Franks, 2021). When there are positivity violations, inverse probability weighting (IPW) estimators (Rosenbaum and Rubin, 1983; Hahn, 1998; Robins et al., 2000; Hirano and Imbens, 2001; Hirano et al., 2003; Imbens, 2004) have substantial statistical bias and inflated variance. A proposed alternative to IPW, direct balancing weights, (Hainmueller, 2012; Zubizarreta, 2015; Wang and Zubizarreta, 2020) tend to yield reduced estimator error and variance compared to IPW methods through achieving exact covariate balance for a prespecified set of covariates. However, when there are positivity violations there is often no solution to the direct balancing optimization problem prohibiting their use altogether.
@@
To mitigate the impacts of positivity violations on causal effect estimation, a variety of methods, such as propensity score trimming, matching weights, and overlap weights, have been developed that modify the estimand and corresponding target population to achieve estimators with reduced variance (Crump et al., 2009; Li and Greene, 2013; Yang and Ding, 2018; Li et al., 2018, 2019). In fact, overlap weights (Li et al., 2018) do not require positivity by targeting the so-called “overlap population”; this results in an estimator with minimum asymptotic variance among weighted treatment effects under homoscedasticity. While these methods achieve reduced estimator variance, they target a population that is modified from the original population of interest such that the corresponding treatment effect estimate may be misaligned with the original research question. Yet, a key benefit of causal effect estimation with observational data is the ability to evaluate causal effects in a meaningful population of interest such that modifying the estimand is often either undesirable or challenging to interpret. Furthermore, these estimators are inconsistent for the original estimand when there is treatment effect heterogeneity. Modified treatment policy estimands result in improved estimator performance by defining an entirely different estimand arrived at by imagining modifications to the natural value of treatment (Kennedy, 2019). While the corresponding estimators perform well, these modified treatment policies do not always answer the original research question.
@@
Thus, under positivity violations, existing methods either 1) target an estimand of interest but result in poor estimator performance or technical issues (e.g., direct balancing weights yielding no solution); or 2) alleviate these technical issues by targeting a modified estimand that may be misaligned with the original estimand of interest (e.g., overlap weights). To address these challenges within estimator statistical performance and interpretation, we propose a novel balancing approach whose optimization problem yields a solution under positivity violations through a targeted relaxation of the balancing constraints. Our corresponding weighted estimator is consistent for the original estimand when either the implied propensity score model is correct or set of treatment effect modifiers is properly specified and smaller than the set of confounders. When these conditions do not hold, our weighted estimator is consistent for a slightly modified estimand and target population. Furthermore, we demonstrate that overlap weights are a special case of our proposed weights such that our method offers an intermediate between direct balancing and overlap weights. We derive asymptotic results which show that 1) our weights can be derived under a relaxed positivity assumption; and 2) our proposed estimator has reduced asymptotic variance when there are positivity violations. We propose a design- and model-based procedure for selecting the balancing constraints to relax and we evaluate the corresponding estimators through applications to synthetic data, an electronic health record (EHR) study, and when transporting a treatment effect from a randomized controlled trial (RCT) to a new population.
@@
The remainder of the paper is organized as follows. In Section 2, we introduce notation and methodological background on balancing weights. We introduce and discuss the implications of our proposed partially retargeted balancing weights in terms of positivity, estimator interpretation, and extensions to additional estimation problems in Section 3. In Section 4, we discuss the asymptotic properties of our proposed weighted estimator. We propose design- and model-based procedures for identifying which balancing constraints to relax in Section 5. In Sections 6 and 7, we apply our methods to synthetic data, an EHR study, and when transporting a treatment effect from a RCT. We discuss results and future work in Section 8.
@@
Consider an independent sample $\{(Z_{i},Y_{i},\bm{X}_{i})\}_{i=1}^{n}$ of size $n$ from a population where $Z_{i}=z$, indicates belonging to the treatment ($z=1$) or control group ($z=0$), $Y_{i}$ is the outcome of interest, and $\bm{X}_{i}=(X_{i1},\ldots X_{ip})$ is a $p$ length vector of pre-treatment covariates. Let $Y(z)$ be the potential outcome that would be observed if assigned to treatment group $z$ (Neyman (1990); Rubin (1974, 1978); Hernan and Robins (2024)). We assume the stable unit treatment value assumption (SUTVA) which implies that $Y_{i}=Y_{i}(Z_{i})$ such that we only observe one potential outcome for each individual. We focus on the average treatment effect (ATE), $\tau=E[Y(1)-Y(0)]$ as the estimand of interest, although this work can be readily extended to other estimands of interest which we discuss in Section 3.3. Let $e(\bm{x})=\Pr(Z_{i}=1|\bm{X}_{i}=\bm{x})$ be the propensity score. For the identification of $\tau$ the following assumptions are typically made:
@@
The focus of our work surrounds the positivity assumption, specifically, scenarios where this assumption does not hold. Let $\mu_{z}(\bm{x})=E[Y(z)|\bm{X}=\bm{x}]$ and $\tau(\bm{x})=\mu_{1}(\bm{x})-\mu_{0}(\bm{x})$. Then, $\tau=E\left\{\frac{\mu_{1}(\bm{X})Z}{e(\bm{X})}-\frac{\mu_{0}(\bm{X})(1-Z)}{1-e(\bm{X})}\right\}=E\left\{\frac{YZ}{e(\bm{X})}-\frac{Y(1-Z)}{1-e(\bm{X})}\right\}$ holds with the above assumptions, where the sample version of the last equation corresponds to the IPW estimator of the ATE. For a treatment effect estimator be consistent, the expectations of $\mu_{0}(\bm{x})$ and $\mu_{1}(\bm{x})$ need to be balanced between the treated and control populations and to the target population. IPW (i.e., $w_{1}(\bm{x})=\{e(\bm{x})\}^{-1}$ and $w_{0}(\bm{x})=\{1-e(\bm{x})\}^{-1}$) balance all functions of $\bm{x}$ in this way such that the following equations hold,
$E\left\{w_{1}(\bm{X})Zb(\bm{X})\right\}=E\{b(\bm{X})\}\>\>$ $\text{and}\>\>E\left\{w_{0}(\bm{X})(1-Z)b(\bm{X})\right\}=E\{b(\bm{X})\}$ (1)
$E\left\{w_{1}(\bm{X})Zb(\bm{X})\right\}$ $=E\left\{w_{0}(\bm{X})(1-Z)b(\bm{X})\right\}$ (2)
@@
for any function $b(\bm{x})$, which we reference as three-way balance (Chan et al., 2016). Thus, IPW satisfy three-way balance of both $\mu_{0}(\bm{x})$ and $\mu_{1}(\bm{x})$, regardless of their true forms. Specifically, Equation (1) balances the weighted treated and control group expectations of $b(\bm{x})$ to the population expectation of $b(\bm{x})$ and Equation (2) balances the weighted treated and control group expectations of $b(\bm{x})$ to each other. Thus, Equation (1) and (2) serve different purposes; while Equation (1) implies Equation (2), Equation (1) primarily serves to balance the treated and control groups to the estimand target population. When only Equation (2) holds, $b(\bm{x})$ may be balanced to modified target population that corresponds not to the ATE, but to some other causal estimand.
@@
However, when there is a lack of overlap, even approximate balance of the empirical counterparts of Equations (1) and (2) may not be achieved after IPW, resulting in estimators with substantial error in addition to inflated variance. To achieve balance when there is a lack overlap, some methods target a different estimand, $E[h(\bm{X})\tau(\bm{X})]$, which corresponds to a modified target population with improved overlap. Li et al. (2018) propose one such method, overlap weights ($h(\bm{x})=e(\bm{x})\{1-e(\bm{x})\}$), which are designed to ensure that the weights exist even when positivity does not hold. By modifying the target population, overlap weights can more easily satisfy Equation (1) yielding an estimator with reduced bias and variance with respect to the modified estimand, the ATO. However, the overlap weights estimator tends to be biased for the ATE when there is treatment effect heterogeneity. While the weights corresponding to these modified estimands tend to achieve improved finite-sample balance in scenarios with a lack of overlap, they rarely achieve the empirical versions of Equations (1) and (2) resulting in estimator error. In fact, even when positivity does hold, exact finite-sample balance is rarely achieved by IPW-style estimators. While exact balance of covariate sample means is achieved by overlap weights when the propensity score is estimated with a logistic regression (Li et al., 2018), most IPW-style weights do not achieve exact balance of the sample means, especially under positivity violations. This has motivated the development of sample weights that explicitly achieve three-way finite sample balance.
@@
$\frac{1}{n}\sum_{i=1}^{n}w_{i}I[Z_{i}=z]b_{j}(\bm{X}_{i})-\frac{1}{n}\sum_{i=1}^{n}b_{j}(\bm{X}_{i})$ $=0\>\>\>j=1,\ldots,J,\>\>\>z=0,1,$ (3b)
where $D(\cdot)$ is a convex measure of dispersion. The constraints $w_{i}\geq 0$ and $\sum_{i=1}^{n}w_{i}Z_{i}=\sum_{i=1}^{n}w_{i}(1-Z_{i})$ are often added to ensure that the weights do not extrapolate the observed outcomes. The resulting direct balancing weights estimator is,
@@
In this constrained optimization problem, constraint (3b) ensures that the treated and control covariate functions are balanced to the sample population. These constraints also imply that the treated and control covariate functions are balanced to each other (i.e., $\frac{1}{n}\sum_{i=1}^{n}w_{i}Z_{i}b_{j}(\bm{X}_{i})=\frac{1}{n}\sum_{i=1}^{n}w_{i}(1-Z_{i})b_{j}(\bm{X}_{i})$) such that three-way balance is satisfied by $\{w^{bw}_{i}\}_{i=1}^{n}$ for $b(\bm{x})$. Since balance between treated and control groups is implicitly satisfied, Problem (3) is separable by treatment group such that two separate constrained optimization problems can be solved to derive $\{w^{bw}_{i}\}_{i=1}^{n}$ (Chan et al., 2016); we discuss implications of this separation in Section 3.1. When there is a lack of overlap between $\{b(\bm{x})\}_{z=1}$ and $\{b(\bm{x})\}_{z=0}$, often there is no solution to Problem (3) (i.e., Equations (3a)-(3b)) with the addition of the constraints $w_{i}\geq 0$ and $\sum_{i=1}^{n}w_{i}Z_{i}=\sum_{i=1}^{n}w_{i}(1-Z_{i})$ such that weights which achieve exact three-way balance without extrapolating cannot be derived. In these scenarios, the exact balance constraint is relaxed and approximate balance is achieved by constraining the absolute difference between sample means to be less than or equal to some parameter $\delta$; weights derived through this process are called minimal weights (Wang and Zubizarreta, 2020). While the minimal weights estimator does have reduced variance compared to IPW and direct balancing estimators, minimal weights achieve none of the three balancing equations such that the corresponding estimator is often biased due to covariate imbalance. Thus, there is a need for methods that 1) achieve exact finite-sample covariate balance between treated and control groups to reduce estimator error; and 2) modify the target population minimally to yield an estimator with reduced variance when there is a lack of overlap.
A partially retargeted balancing approach
To address these challenges, we propose a novel constrained optimization problem for deriving weights, primarily for scenarios where there is no solution to direct balancing weights due to positivity violations. Consider the case where Assumption Assumption 2: Positivity does not hold, but $0<\text{Pr}\{Z_{i}=1|c(\bm{X})=c(\bm{x})\}<1$ for some $c(\bm{x})\subset b(\bm{x})$ such that positivity holds when conditioning on this restricted subset. Let $g(\bm{x})$ be the remaining covariate functions in $b(\bm{x})$ such that $b(\bm{x})=[c(\bm{x})=\{c_{k}(\bm{x})\}_{k=1}^{K},g(\bm{x})=\{g_{l}(\bm{x})\}_{l=1}^{L}]$. The key idea of our method is to relax the balancing constraints by balancing all covariate functions, $b(\bm{x})=\{c(\bm{x}),g(\bm{x})\}$, between treated and control groups, but only $c(\bm{x})$ to the sample population, yielding the following set of balancing equations:
@@
The primary difference between this constrained optimization problem and the direct balancing weights constrained optimization problem is Equation (5c); in this equation, the functions, $g(\bm{x})=\{g_{l}(\bm{x})\}_{l=1}^{L}$ are only balanced between treated and control groups and not to the sample population. Thus, the derived weights satisfy exact three-way balance for functions $c(\bm{x})=\{c_{k}(\bm{x})\}_{k=1}^{K}$ but only balance between treated and control groups for $g(\bm{x})$. In doing so, the derived weights do not balance $g(\bm{x})$ to a specific modified population (such as with overlap weights). However, the weighted treated and control group sample means of $g(\bm{x})$ are implicitly modified when minimizing weight dispersion subject to only the treated and control balance constraint. We call these weights, $\{w_{i}^{*}\}_{i=1}^{n}$, partially retargeted balancing weights (PRTBW). In Section 3.1, we formally show that this approach is justified under a relaxed positivity assumption. In addition, the corresponding estimator to Problem (5) (i.e., Equations (5a) - (5c)) is consistent for the ATE when $c(\bm{x})$ contains all treatment effect modifiers, which we discuss further in Sections 3.2 and 4.
@@
Figure 1: Simulated data illustrating the differences between direct balancing weights (BW), minimal weights (MW), and the proposed partially retargeted balancing weights (PRTBW). Squares (purple) and triangles (orange) indicate the unweighted control and treated sample means, respectively. The filled diamond indicates the sample population mean while the outlined diamond indicates the weighted sample mean given $\{w_{i}^{*}\}_{i=1}^{n}$ derived through Problem (5). In the right-hand figures, the horizontal line indicates the sample mean while the vertical lines end at the weighted treated and control means for each weighting method. For this simulated data, there is no solution to Problem (3), exhibited by the dashed vertical line.
@@
Figure 1 presents a toy example of the differences in covariate balance and target population between balancing weights (Problem (3)), minimal weights, and our proposed partially retargeted balancing weights (Problem (5)); for this example, the constraints of the $w_{i}\geq 0$ and $\sum_{i=1}^{n}w_{i}Z_{i}=\sum_{i=1}^{n}w_{i}(1-Z_{i})$ are added to the optimization problems. For this simulated data, there is no solution to Problem (3), such that weights that exactly balance the treated and control sample means of $b(\bm{x})=\{c(\bm{x}),g(\bm{x})\}$ cannot be derived. Our proposed weights balance the treated and control sample means of $c(\bm{x})$ to the population sample mean, but the treated and control sample means of $g(\bm{x})$ are balanced to a different value. In contrast, minimal weights do not achieve covariate balance for either $c(\bm{x})$ or $g(\bm{x})$. In fact, the minimally weighted treated sample means are unchanged and the minimally weighted control sample mean of $c(\bm{x})$ is further away from the population sample mean of $c(\bm{x})$ than the unweighted control sample mean. Thus, while both the minimal weights optimization problem and our proposed Problem (5) yield solutions for this data, they do this through different mechanisms such that each method yields a solution with a different interpretation.